8. A credit card company charges 15.6% percent per year interest. Compute the effective annual rate if they compound, (a) annualy, (b) monthly (c) daily and (d) continuosuly.
Answer #1

8)

Effective annual rate =  (1+APR/n)n -1

a) Annualy

that is n = 1

EAR = (1 + 0.156)1 -1

EAR = 1.156 -1 = 0.156 = 15.6%

b) Monthly means n = 12

EAR = (1+0.156/12)12 -1

= 1.01312 -1

= 1.167 -1

= 0.167 = 16.7%

c) daily

n = 365

EAR = ( 1 + 0.156/365)365 - 1

= (1.000427)365 -1

= 1.1686 - 1

= 0.1686 = 16.8%

Hope this helps

please comment if you have any doubts.